You signed a promissory note with a bank that requires you to pay $4000 in 5 years. If the money is receiving an interest rate of 9% compounded quarterly, what amount must be paid today to discharge (pay off) the debt early?
Answer by Theo(13342)

that would be the present value of 4000 at (9/4)% per quarter for 5 * 4 = 20 quarters.
your formula would be:

f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

the time periods are quarters of a year.
5 years * 4 quarters per year = 20 quarters.

the interest rate is the annual interest rate divided by 4 = 9/4 = 2.25%.
in the formula, the rate is used, not the percent, so the interest rate per quarter = .0225.

since the future value is 4000, the formula becomes:

4000 = p * (1 + .0225) ^ 20

solve for p to get:

p = 4000 / (1 + .0225) ^ 20 = 2563.265887.